My Notes o Ask Your T e 46 points Previous Answers CJ9 18 P014 Go. The drawings show three charges that have the same magnitude but may have different signs. In all cases the distance d between the charges is the same. magnitud of the charges is lal- 2.3 pc, and the distance between them is d 3.5 mm. Determine the magnitude of the net force on charge 2 for each of the three drawings. (a) F (b) F (c) F
Answer #1

part a:

force between 1 and 2 is attractive and hence along -ve x axis.

force between 2 and 3 is repulsive and hence along -ve x axis direction.

so both the forces are additive.

magnitude of each of the forces are equal

magnitude of each force=9*10^9*2.3*10^(-6)*2.3*10^(-6)/(3.5*0.001)^2

=3886.53 N

so total force=2*3886.53=7773 N

part b:

due to symmetry, total force =0

part c:

force between 1 and 2 is attractive and hence along -ve x axis.

force between 2 and 3 is attractive and hence along -ve y axis direction.

magnitude of each of the forces are equal

magnitude of each force=9*10^9*2.3*10^(-6)*2.3*10^(-6)/(3.5*0.001)^2

=3886.53 N

magnitude of total force=sqrt(force along x axis^2+force along y axis^2)

=5496.383 N